# Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives

### v1

@@ 2.3.3 Micro Level
As AI agents become increasingly embedded in organizational operations and business processes, the issue of value alignment at the micro level has grown more prominent. Unlike the overarching constraints imposed by macro-level universal values and meso-level policy regulations, the micro level focuses on how AI systems achieve alignment between effectiveness and values within localized environments under specific goal orientations. The scenarios and associated values at this level are inherently more personalized. Given the impossibility of exhaustively enumerating all possible scenarios, this study synthesizes relevant research and illustrates the principles of micro-level value alignment through examples such as talent recruitment, legal consultation and corporate governance in the healthcare sector.
@@
The application of artificial intelligence technologies is profoundly reshaping human resource management practices, with the recruitment domain emerging as a primary frontier of technological transformation. AI-powered talent selection systems are driving a fundamental shift in organizational talent acquisition paradigms. Compared to traditional recruitment models that rely heavily on manual screening and subjective judgment, intelligent tools significantly enhance the speed and accuracy of candidate identification through automated resume parsing, big data-driven competency modeling, and machine learning-enabled predictive assessments [97].The ethical compatibility of AI technologies in recruitment settings has attracted sustained academic attention. In constructing ethical principles, Simbeck advocates for the transfer of key moral concepts from other domains to human resource analytics when applying new AI technologies, identifying five core principles: data privacy protection, user opt-out mechanisms, third-party audit frameworks, algorithmic transparency, and the safeguarding of dynamic development rights [79]. These principles align with the humanistic view of technology proposed by Runb-Kettler and Lehnervp, which emphasizes preserving candidates’ dignity and autonomy in automated processes [78]. A systematic literature review by Hunkenschroer and Luetge further deconstructs the ethical dimensions of AI in recruitment into ethical opportunities, ethical risks, and areas of ambiguity [80].
From the perspective of ethical opportunities, AI can reduce human biases in traditional recruitment through standardized assessment procedures [98][99]. Text optimization tools can help generate gender-neutral job descriptions [100], while natural language processing has revolutionized resume parsing efficiency, allowing recruiters to focus on strategic decision-making [101]. However, ethical risks remain significant. Algorithms may inherit implicit biases embedded in training data [102], implement indirect discrimination via proxy variables (e.g., zip codes correlated with race), or infringe on privacy rights by extracting personal information from candidates for recruitment purposes [103][104]. Although AI has the potential to disrupt traditional elite hiring models and promote social mobility, it may also lead to reduced cognitive diversity in teams due to algorithmic homogenization. These tensions and risks underscore fairness as a core principle in the application of AI to human resource recruitment.
@@
The deep integration of AI technologies is also influencing the strategic development and governance choices of biopharmaceutical companies. In response to the ethical challenges posed by the application of AI in healthcare, AstraZeneca has issued a series of AI ethical principles, including privacy and security, interpretability and transparency, fairness, accountability, human-centeredness, and societal benefit. The company has also developed a Responsible AI Handbook, established an AI Resolution Committee and internal advisory services, and commissioned AI audits to implement comprehensive AI governance [82].
In addition, AI technologies are demonstrating significant potential and application prospects in the legal consulting domain. Sun et al. simulated the collaborative workflows of real-world law firms using multi-agent systems. In this context, they incorporated normative principles—such as privacy protection, fairness, transparency, and accountability—to guide the ethical behavior and interactions of the agents. This alignment with human values enhances the quality of legal consulting services by ensuring that multi-agent systems operate in a manner consistent with ethical and societal expectations [81].
@@
3 Applications and Value Issues in LLM-based agent systems
@@
LLM-based agent systems are now applied across numerous domains, indicating that value alignment concerns will permeate to the application scenario level. This section first introduces our review subjects—LLM-based agents and multi-agent systems. Subsequently, it classifies and organizes intelligent agent system application scenarios based on their generality, and discusses the associated value alignment issues.
3.1 LLM-Based Agent Systems
3.1.1 LLM-Based Agent
In philosophy, the term ’agent’ typically refers to an entity possessing consciousness, intention, and the capacity for action [105, 106]. This can include humans, animals, and other entities in both the physical and virtual worlds [1]. Floridi identifies three general characteristics of an agent: interactivity, autonomy, and adaptivity. Interactivity refers to an agent’s ability to interact with and respond to environmental stimuli. Autonomy indicates an agent’s capacity to change its own state based on its transformation rules. Adaptivity denotes an agent’s ability to not only alter its state but also to modify the rules governing these changes through learning and experience, thereby flexibly responding to environmental shifts. Building on these three characteristics, Floridi posits that a Moral Agent can exhibit morally evaluable behavior, meaning their actions can engender moral "good" or "evil." Consequently, Moral Agents must adhere to specific ethical norms and oversight mechanisms [107]. When the concept of an "Agent" is introduced into the field of artificial intelligence, it becomes known as an AI agent. An AI agent is an artificial entity capable of achieving complex goals through collaboration [108], representing the concretization of the philosophical concept of "agent" within artificial intelligence.
Large Language Models (LLMs) are deep learning models trained on vast amounts of textual data. They are capable of generating natural language text and comprehending its meaning, with examples including Deepseek, GPT-3 and GPT-4. A salient characteristic of LLMs is their adherence to scaling laws and the exhibition of emergent abilities. This means that once a model’s training parameters exceed a certain threshold, it can experience a dramatic improvement in existing capabilities or acquire entirely new ones [17], enabling it to solve tasks it was not explicitly trained for [12]. As models continue to grow in scale, pre-training corpora expand, and alignment techniques advance, LLMs have demonstrated powerful capabilities in knowledge acquisition, instruction understanding, summarization, planning, and reasoning [1]. This showcases their immense potential for achieving human-like intelligence. Consequently, researchers have begun leveraging LLMs to construct intelligent digital agents [23], transforming abstract philosophical concepts into practically engineerable system design frameworks [1].
AI Agents are intelligent entities capable of perceiving their environment, making decisions, and executing actions [109, 110, 111]. LLMs can serve as the central controller for AI agents, endowing them with decision-making thought and memory [112]. Decision-making thought enables agents to decompose complex tasks, deliberate on each sub-task, and explore multiple pathways to arrive at intricate decisions. A long- and short-term memory mechanism allows agents to maintain contextual coherence during task completion and continuously learn from past interactions. Strategies such as multimodal perception and tool use expand an agent’s perception and action space [113, 114, 115, 116, 117]. Furthermore, AI agents endowed with environmental interaction capabilities can learn from feedback and execute novel actions [118]. The development of techniques like Chain-of-Thought (CoT) and problem decomposition has continuously enhanced the reasoning and planning capabilities of AI agents[119, 120]. Furthermore, pre-training on massive-scale corpora has enabled agents to exhibit few-shot and zero-shot generalization abilities. This allows them to transfer and complete tasks across different scenarios without requiring parameter updates [120, 121]. The field of AI is currently undergoing a profound paradigm shift. Building upon existing research, the industry is transitioning from developing AI models for passive, structured tasks to constructing novel models capable of acting as dynamic agents in diverse and complex scenarios.
3.1.2 Multi-Agent System
A Multi-Agent System (MAS) is a computer system composed of multiple interacting agents [12]. Operating on natural language and custom content [24], these agents can interact within a shared environment, autonomously making decisions to solve complex problems or complete tasks [122]. The key components of such a system include the Agent, Environment, Interaction rules, and Organization pattern [12].
  • Agent.The agent is the core participant in the system, possessing a role identity, knowledge, capabilities, and behaviors, along with abilities such as learning, planning, reasoning, and decision-making.
  • Environment.The environment is the external world in which the agent operates, which can be either a simulated virtual environment or a real physical space.
  • The Interaction Rules. The rules are predefined rules, protocols, and mechanisms governing information exchange and behavior coordination among agents.
  • The Organization Pattern. The pattern refers to the structural relationships of agent interactions, encompassing hierarchical control and emergent behavior-based organizations.
Multi-agent interaction within these systems mirrors the specialized division of labor and teamwork found in human society; this aims to achieve collective intelligence. Individual autonomous agents possess initial knowledge and goal sets tailored for specific tasks [122]. By leveraging LLMs’ capabilities in text generation, comprehension, and communication, diverse agents can adaptively collaborate. They respond to task requirements, participate in planning, discussion, and decision-making, and learn from one another. This fosters the aggregation of broad knowledge across multiple domains [112, 12]. Concurrently, through environmental interaction, agents perceive and understand information. They then make decisions and execute actions based on their goals, existing knowledge, and observations. This process unleashes their potential to complete specialized tasks and achieve common objectives.
Multi-agent systems exhibit remarkable flexibility, enhanced robustness, inherent self-organization and coordination, and robust real-time operation capabilities [12]. Although interactions among agents are governed and constrained by predefined rules, emergent behaviors and phenomena can still arise from these interactions [123]. Additionally, due to the division of labor within the system, resource and energy consumption are distributed across multiple agents. This is often more cost-effective than a single, powerful agent attempting to solve an entire complex problem [12]. Therefore, compared to a single agent, multi-agent systems are better suited for handling more complex and dynamic tasks, particularly in scenarios requiring diverse strategies and collective decision-making [112]. At present, LLM-based multi-agent systems are extensively used to simulate social phenomena. This includes modeling social network construction, international conflicts, political evolution, and the dynamics of public opinion  [124, 125, 126].

### v2

@@ 2.4.3 Micro Level
As AI agents become increasingly embedded in organizational operations and business processes, the issue of value alignment at the micro level has grown more prominent. Unlike the overarching constraints imposed by macro-level universal values and meso-level policy regulations, the micro level focuses on how AI systems achieve alignment between effectiveness and values within localized environments under specific goal orientations. The scenarios and associated values at this level are inherently more personalized. This study synthesizes relevant research and illustrates the principles of micro-level value alignment through examples such as talent recruitment, legal consultation and corporate governance in the healthcare sector.
@@
AI technology is fundamentally reshaping human resource management practices, with recruitment being the forefront of this transformation. AI-based talent selection systems are revolutionizing organizational talent acquisition, improving speed and accuracy, compared to traditional manual screening and subjective judgment [86]. The ethical adaptation of AI in recruitment has attracted growing academic attention. In terms of ethical principles, Simbeck proposed five core principles for applying AI in human resource analytics: data privacy, user exit mechanisms, third-party review, algorithm transparency, and the right to dynamic development [72]. These principles align with Rünb-Kettler and Lehnervp’s humanistic technological view, which emphasizes maintaining candidate dignity and subjectivity in automated processes [71].
From an ethical opportunity perspective, AI reduces human bias in traditional recruitment [87][88], generates gender-neutral job descriptions [89], and revolutionizes resume parsing efficiency with natural language processing, allowing recruiters to focus on strategic decisions [90]. However, ethical risks remain, such as algorithms inheriting hidden biases from training data [91], implementing indirect discrimination through proxy variables (e.g., postal code linked to race), and privacy violations through the extraction of personal information for recruitment purposes [92][93]. These contradictions and risks make fairness the core principle in AI recruitment applications.
@@
AI technology’s deep integration is also influencing strategic development and governance choices in biopharmaceutical companies. To address the ethical challenges posed by AI in healthcare, AstraZeneca has introduced a series of AI ethics principles, including privacy and security, explainability and transparency, fairness, accountability, human-centeredness, and social benefit [75].
Additionally, AI technology shows potential and application prospects in the legal consulting field. Sun et al. simulated multi-agent collaboration in a real law firm setting and incorporated privacy, fairness, transparency, and accountability principles to guide multi-agent ethical behavior and interaction, aligning them with human values and enhancing the quality of legal consulting services [74].
@@
3 Applications and Value Issues in Agentic AI systems
@@
Agentic AI systems are increasingly deployed across diverse domains, indicating that value alignment concerns will extend to the application scenario level. This section classifies the application scenarios of agentic AI systems based on their generality and discusses the associated and important value alignment issues.